A typical gas cylinder used for such depths contain 69.33g of O2 and 457.24g of N2 and has a volume of 20.00L. What is the partial pressure of each gas at 15 degrees, and what is the total pressure in the cylinder at this temperature?

MM(O2)= 32.00 u
MM(N2)=28.0u
R= 0.0821L

Partial pressure of O2=
Partial pressure of N2=
Total pressure in the tank=
1
Expert's answer
2020-11-22T11:14:32-0500

In calculating the number of moles of nitrogen and oxygen gas, we use the ideal gas law;

"n=" "PV\over RT"

"nO_2=" "(1atm)(0.06933L)\over (0.0821{atm.L\over mol.K})(273+15k)" "=2.9\times 10^{-3} mol O_2"

"nN_2=" "(1atm)(0.45724L)\over (0.0821L{atm.L\over mol.L})(273+15k)"

"=1.9\times 10^{-2}mol N_2"

In calculating the partial pressure of the gases, we use Daltons equation

"P=" "nRT\over V"

"P_{O_2}=" "(2.9\times 10^{-3})(0.0821{atm.L\over mol.L})(288k)"

"=0.0685atmO_2"

"P_{N_2}=(1.9\times 10^{-2})(0.0821{atm.L\over mol.L})(288k)"

"=0.449atm N_2"

"P_{Total}=P_{N_2}+P_{O_2}"

"=0.449atm+0.0685atm"

"P_{Total}=0.5175atm"